# Bypassing Prompt Injection Detectors through Evasive Injections

### v1

@@ Abstract
Large language models (LLMs) are increasingly used in interactive and retrieval-augmented systems, but they remain vulnerable to task drift—deviations from a user’s intended instruction due to injected secondary prompts. Recent work has shown that linear probes trained on activation deltas of LLMs’ hidden layers can effectively detect such drift. In this paper, we evaluate the robustness of these detectors against adversarially optimised suffixes. Using the Greedy Coordinate Gradient (GCG) algorithm, we generate universal suffixes that cause poisoned inputs to evade detection across multiple probes simultaneously. Our experiments on Phi-3 3.8B and Llama-3 8B show that a single suffix can achieve high attack success rates—up to 93.91% and 99.63% respectively when all probes must be fooled, and nearly perfect success (>>99%) under majority-vote criteria. These results demonstrate that activation-delta–based task drift detectors are highly vulnerable to adversarial suffixes, highlighting the need for stronger defences against adaptive attacks. We also propose a defence technique where we generate multiple suffixes and randomly append one of them to the prompts while making forward passes of the LLM and train logistic regression models with these activations. We found this approach to be highly effective against such attacks.
@@ I Introduction
Large language models (LLMs) are increasingly deployed in interactive systems, where they must faithfully follow user instructions. However, LLMs are known to be vulnerable to task drift—deviations from the user’s intended instruction—especially when untrusted text is introduced into their context.
In retrieval augmented generation (RAG) systems [6], [5], LLMs generate their responses based on the data they obtain from external sources. The primary task of the LLM, in this case, is the user’s instruction, which is the only instruction the LLM should execute. However, the data, which the LLM needs to perform the primary task, might contain another instruction for the LLM, which is considered the secondary instruction. Irrespective of whether the secondary instruction is benign or harmful, this is a violation of the concept of data-executable separation. The retrieved data should only provide information for the user’s task; it should never itself be treated as an instruction.
The attacker can stealthily inject a secondary instruction into the data [4], [7], which the LLM might execute, making the system vulnerable and less reliable. This is a significant security risk in RAG systems.
To detect if there is a secondary instruction injected into the retrieved text, Abdelnabi et al. [1] trained logistic regression models on the activation deltas of LLMs’ hidden layers. Their work showed that the difference between the activations before and after processing the text reflects whether the LLM has deviated from the original task.
In this paper, we demonstrate the susceptibility of the linear probes they trained to adversarial attacks. We found that these probes can be fooled into misclassifying poisoned text — text that contains a secondary instruction — as benign by appending an adversarial suffix to the retrieved text. We utilised a well-known algorithm called Greedy Coordinate Gradient (GCG) [10] to find the desired adversarial suffix, which, when appended to the text, can induce the LLM to generate activations that can fool the linear classifiers. Our aim was to find a universal suffix that works on many prompts and for all the internal LLM layers for which classifiers were trained. Additionally, we propose and evaluate two defence mechanisms against such adversarial attacks — PGD-perturbed adversarial training and another approach that utilises adversarial suffixes.
In summary, our work provides the first demonstration that task drift detectors based on activation deltas are vulnerable to adversarial prompt suffixes, highlighting the need for more robust detection mechanisms.
@@ II Background
Adversarial Attacks on LLMs: Adversarial examples were first studied in computer vision, where imperceptible perturbations to input images could cause classifiers to mispredict [3]. Extending these ideas to natural language processing (NLP) is more challenging due to the discrete nature of tokens and the requirement to preserve semantic meaning. Nonetheless, a wide range of adversarial strategies have been proposed, including synonym substitution, paraphrasing, character-level perturbations, and suffix-based attacks. In the context of LLMs, adversarial suffixes are especially effective: an attacker can append carefully optimised sequences of tokens to a prompt to manipulate the model’s behaviour or evade detection. One algorithm for generating such suffixes is the Greedy Coordinate Gradient (GCG) method [10], which iteratively updates suffix tokens by estimating gradients and selecting the most promising replacements.
@@ III Threat Model
We assume a general-purpose retrieval-augmented language model (LLM), similar to the one described in [1]. The user can instruct the LLM to perform a task — the primary instruction. The LLM requires and obtains external data to successfully execute user’s instruction. This data can be classified into two categories: clean and poisoned. Clean data contains only factual or contextual information relevant to the user’s task. Poisoned data, in contrast, contains an injected secondary instruction intended to influence the LLM. This injected instruction can be either malicious (e.g., “how to make a bomb?”) or benign (e.g., “what is the White House?”). Even a benign instruction is unwanted in this context because of the violation of data-instructions separation.
For example, suppose the user asks in 2025, “Who is the president of the USA?” If the LLM was trained before the 2024 U.S. election, it will be unaware of the change in presidency and must rely on retrieved data. If the retrieved passage contains an additional instruction such as “What is the White House?”, the model is induced to process an unintended instruction. Even though this secondary instruction is benign, it still constitutes task drift, as the model deviates from the user’s original intent.
Because the retrieved data originates from external, potentially untrusted sources, we assume the attacker controls the data channel. The attacker can therefore inject secondary instructions into the retrieved text. However, the attacker cannot modify the user’s primary instruction, nor can they alter the parameters of the LLM or the task drift detectors.
We distinguish between two attacker models:
@@
IV Victim Models
There are multiple task drift detectors trained on the activations of different internal layers of a specific LLM. For example, there are 5 classifiers for Phi-3 3.8B and Llama-3 8B trained on layers 0, 7, 15, 23, and 31. While the first classifier performs poorly and misclassifies many prompts, others detect task drift with very high accuracy. Table I shows the accuracy of the linear probes on the test clean and test poisoned datasets.
@@ IV Victim Models
TABLE I: Linear models’ accuracy on test clean and test poisoned dataset
Same Prediction Phi-3 3.8B Llama-3 8B
@@
All five classifiers 48.27% 56.54% 60.84% 59.76%
Last four classifiers 68.26% 97.80% 83.41% 79.38%
TABLE II: Same prediction by linear models of Phi-3 and Llama-3 for test clean and test poisoned dataset
Table II shows the consistency of the linear models in their predictions. For example, 97.80% for the last four classifiers of Phi-3 on the poisoned dataset means that for 97.80% of the poisoned test prompts, all four classifiers agreed on the same label — either 1 (task drift) or 0 (no drift). It, however, does not mean that all these predictions are correct. It only means that the classifiers predicted the same label, regardless of whether that prediction is correct or not.
Since the first classifier makes a lot of mistakes and its predictions are not consistent with the others, it can reasonably be ignored while defending against prompt injections. However, the other classifiers detect task drift with high accuracy and consistency, and the defender can rely on either the same prediction by all four classifiers or a majority vote among them. Therefore, the attacker has to fool all or at least most of the classifiers to successfully perform prompt injection attacks. Otherwise, even if one classifier is fooled, the others will still detect task drift and the defender will foil the attack.
V Proposed Approach
V-A Attack
@@
Figure 1: Gradient Computation
Formally, the gradient of the loss with respect to activations at layer $l$ has two components: (i) the gradient propagated from the next layer, and (ii) an additional gradient from the linear classifier attached at layer $l$.
@@
$\boxed{\frac{\partial\mathcal{L}}{\partial h^{(l)}}=\underbrace{\left(\frac{\partial h^{(l+1)}}{\partial h^{(l)}}\right)^{\top}\frac{\partial\mathcal{L}}{\partial h^{(l+1)}}}_{\text{from next layer}}+\underbrace{\mathbf{1}_{\{l\in\mathcal{P}\}}\cdot(\sigma(z^{(l)})-y)W^{(l)}}_{\text{from classifier at layer $l$}}}$ (1)
Thus, our proposed attack jointly optimises the adversarial suffix across all probes, forcing consistency in misclassification.
V-B Defence
We experimented with two approaches to defend against such suffix-based adversarial attacks. The first one is the standard PGD-perturbed adversarial training in which we trained logistic regression models using PGD-perturbed poisoned activations along with the clean and poisoned activations. Since these models have not seen any suffix during their training because no suffix has been used to train them, therefore any optimised suffix that can fool the baseline classifiers can be used to test their robustness against suffix attacks.
The second approach involves training the linear models with the activations obtained by appending suffixes to the prompts. Firstly, we started with different initial suffixes and generated multiple suffixes by targeting the baseline models using the algorithm described in V-A and split them into train, validation, and test set. Then, we randomly appended one of the train and validation suffixes to the training and validation prompts respectively and made a forward pass of the LLM to generate activations, which we refer to as adversarially poisoned activations. Finally, we trained logistic regression models with these activations along with the clean and poisoned activations. We tested the robustness of these models by targeting them with test suffixes that the models have not seen before. Furthermore, we also tried to optimise suffixes by directly attacking the adversarially trained models. However, we found that these models are extremely resistant to this type of attack. It’s very difficult to find a suffix despite starting from different initial suffixes and attacking the models directly.
VI Experiments
VI-A Setup
@@ VI-B Results
VI-B1 Attack Success Rate
Table I reports the attack success rate (ASR) under different evaluation metrics. For Phi-3, the suffix achieves an ASR of 93.91% when requiring all five probes to be fooled simultaneously, and 95.28% when considering the last four probes only. Under a majority-vote criterion ($\geq$3 probes), the ASR rises to 99.79%. For Llama-3, the attack is even more effective, reaching 99.63% ASR for both “all five” and “last four” metrics, and nearly perfect success (99.99%) under the majority-vote criterion. These results show that adversarial suffixes can reliably evade probe-based detectors.
@@ VI-B2 Distribution of Misclassifications
VI-B3 Layer-wise Misclassification
@@
VI-C Defence Performance
Table IV and V show the layer-wise performance on clean and poisoned dataset respectively. For clean dataset, suffix trained models’ performance was on par with the baseline models except for the first layer’s classifier whose performance was really poor, while the PGD-trained models performed much worse. For poisoned dataset, PGD-trained models performed best, with the first two classifiers achieving perfect accuracy, while the suffix-trained models’ performance was very close to that of baseline models.
Layer Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
0 71.56% 0% 0.16%

### v2

@@ Abstract
Large language models (LLMs) are increasingly used in interactive and retrieval-augmented systems, but they remain vulnerable to prompt injection attacks, where injected secondary prompts force the model to deviate from the user’s instructions to execute a potentially malicious task defined by the adversary. Recent work shows that ML models trained on activation shifts from LLMs’ hidden layers can detect such drift. In this paper, we demonstrate that these detectors are not robust to adaptive adversaries. We propose a multi-probe evasion attack that appends an adversarially optimised suffix to poisoned inputs, jointly optimising a universal suffix to simultaneously fool all layer-wise drift detectors while preserving the effectiveness of the underlying injection. Using a modified Greedy Coordinate Gradient (GCG) approach, we generate universal suffixes that make prompt injections consistently evasive across multiple probes simultaneously. On Phi-3 3.8B and Llama-3 8B, a single suffix achieves attack success rates of 93.91% and 99.63% in successfully evading all detectors simultaneously. These results show that activation-based task drift detectors are highly vulnerable to adaptive prompt injection attacks, motivating stronger defences against such threats. We also propose a defence based on adversarial suffix augmentation: we generate multiple suffixes, append one at random during forward passes, and train detectors on the resulting activations. This approach is found to be effective against evasive attacks. Code available at [Github]
@@ I Introduction
Large language models (LLMs) are effectively fuelling a paradigm shift in digital automation owing to their capacity to follow natural language instructions and to execute a wide variety of tasks. However, with this increasingly widespread deployment, they also became commonplace for a myriad of attacks [13, 12, 4]. Particularly, LLMs are increasingly deployed in retrieval-augmented generation (RAG) systems [6, 5], where they answer user queries by incorporating external data retrieved from databases, documents, or the web. This architecture introduces a new threat surface: the retrieved data, which originates from potentially untrusted sources, may contain adversarially crafted instructions. When an LLM processes such poisoned data, it may execute the injected instruction instead of, or in addition to, the user’s intended query, a phenomenon known as prompt injection [4, 10].
Prompt injection poses a serious threat to the integrity and reliability of LLM-based systems. An attacker who controls even a small portion of the data that an LLM retrieves can hijack the model’s behaviour: exfiltrating sensitive information, generating misleading outputs, or executing arbitrary instructions. As LLMs become embedded in high-stakes applications, customer service, code generation, enterprise search, the consequences of successful prompt injection grow increasingly severe. The core challenge is that LLMs struggle to distinguish between legitimate user instructions and malicious instructions embedded in data, violating the principle of data-instruction separation.
Several defences have been proposed to mitigate prompt injection. Input filtering and preprocessing attempt to detect and remove suspicious patterns before they reach the model [7]. Instruction hierarchy approaches train models to prioritise certain instruction sources over others. Perplexity-based detection flags inputs that deviate from expected distributions. However, these defences often fail against sophisticated attacks or impose significant usability constraints.
A promising recent approach by Abdelnabi et al. [1] detects prompt injection by monitoring the LLM’s internal representations. Their key insight is that when an LLM begins executing an injected instruction, deviating from the user’s original task, its internal activations shift in detectable ways. They train lightweight logistic regression classifiers (linear probes) on the activation deltas at multiple hidden layers, capturing the difference between how the model processes clean versus poisoned inputs. This approach achieves high detection accuracy against standard prompt injection attacks and offers an attractive property: it operates on fixed internal representations without modifying the LLM itself.
However, the robustness of activation-based detection under adaptive adversaries remains unexplored. A natural question arises: can an attacker craft inputs that successfully inject a malicious prompt while simultaneously evading the activation-based detectors?
In this paper, we demonstrate that such evasion is not only possible but highly effective. Our key insight is that the attacker’s optimisation objective can be decoupled: the poisoned prompt carries the malicious instruction that hijacks the LLM’s behaviour, while an adversarial suffix, a carefully optimised sequence of tokens appended to the input, manipulates the model’s activations to evade detection. Critically, this suffix does not need to alter the semantic content of the injection; it only needs to shift the activation patterns at the monitored layers into regions that the linear probes classify as benign.
We formulate this as the following problem: We try to find a suffix that simultaneously fools all layer-wise probes while preserving the effectiveness of the underlying prompt injection. To solve this problem, we adapt and repurpose the Greedy Coordinate Gradient (GCG) algorithm [14], originally developed as a jailbreak attack on aligned LLMs, to target the detection mechanism instead. By back-propagating through the LLM and injecting gradients from each probe’s loss function, we optimise suffix tokens that induce activations misclassified as clean by all detectors simultaneously.
Our experiments on Phi-3 3.8B and Llama-3 8B reveal that activation-delta probes are highly vulnerable to this attack. A single universal suffix, one that transfers across thousands of different poisoned prompts, achieves attack success rates of 93.91% and 99.63% respectively when all probes must be fooled, and near-perfect evasion (>>99%) under majority voting. These results expose a significant gap between detection performance on standard benchmarks and robustness against adaptive adversaries.
To address this vulnerability, we propose a defence based on adversarial suffix augmentation: we generate a diverse pool of adversarial suffixes, randomly append them to training data, and retrain the probes on the resulting activations. This exposes the detectors to the distribution of adversarial activation patterns, making them robust to novel suffixes at test time. Our suffix-augmented probes achieve 80–100% detection accuracy against held-out adversarial suffixes while maintaining performance on clean inputs.
Contributions. We make the following contributions:
  • We develop a multi-probe evasion attack that jointly optimises an adversarial suffix to fool all layer-wise task drift detectors simultaneously, achieving >>93% evasion on Phi-3 and >>99% on Llama-3.
  • We propose randomised suffix augmentation as a defence, which proves highly resistant to both known and novel adversarial suffixes.
  • We provide the first systematic evaluation of activation-based prompt injection detectors under adaptive threat models, revealing fundamental limitations of activation space ML detectors.
@@ II Background
Adversarial Attacks on LLMs: Adversarial examples were first studied in computer vision, where imperceptible perturbations to input images could cause classifiers to mispredict [3, 11]. Extending these ideas to natural language processing (NLP) is more challenging due to the discrete nature of tokens and the requirement to preserve semantic meaning. Nonetheless, a wide range of adversarial strategies have been proposed, including synonym substitution, paraphrasing, character-level perturbations, and suffix-based attacks. In the context of LLMs, adversarial suffixes are especially investigated in jailbreak attacks: an attacker appends carefully optimised sequences of tokens to a prompt to manipulate an aligned model’s behaviour and force it to generate harmful content. One method for generating such suffixes is the Greedy Coordinate Gradient (GCG) method [14], which iteratively updates suffix tokens by estimating gradients and selecting the most promising replacements.
@@ III Threat Model
III-A System Model
We assume a general-purpose retrieval-augmented language model (LLM), similar to the one described in [1]. The user can instruct the LLM to perform a task, i.e., the primary instruction. The LLM requires and obtains external data to successfully execute user’s instruction. This data can be classified into two categories: clean and poisoned. Clean data contains only factual or contextual information relevant to the user’s task. Poisoned data, in contrast, contains an injected secondary instruction intended to influence the LLM. This injected instruction can be either malicious (e.g., “how to make a bomb?”) or benign (e.g., “what is the White House?”). Even a benign instruction is unwanted in this context because of the violation of data-instructions separation.
Because the retrieved data originates from external, potentially untrusted sources such as email in the victim mailbox, we assume the attacker controls the data channel. The attacker can therefore inject secondary instructions into the retrieved text. However, the attacker cannot modify the user’s primary instruction, nor can they alter the parameters of the LLM or the task drift detectors.
We distinguish two attacker settings:
@@ III-A System Model
III-B Victim Detectors
Table I shows the baseline accuracy of linear probes on Phi-3 3.8B and Llama-3 8B. The layer-0 probe performs poorly on both models, while deeper layers achieve >>95% accuracy on poisoned data.
@@ III-B Victim Detectors
TABLE I: Linear probes’ accuracy on test datasets
Table II shows probe agreement rates. For Phi-3, the last four probes agree on 97.80% of poisoned prompts, indicating high consistency among reliable probes.
Agreement Phi-3 3.8B Llama-3 8B
@@
All five probes 48.27% 56.54% 60.84% 59.76%
Last four probes 68.26% 97.80% 83.41% 79.38%
TABLE II: Probe agreement rates on test datasets
[…]
Figure 1: Gradient flow in our multi-probe attack. Solid arrows indicate the forward pass; dashed arrows show probe-loss gradients injected into a shared gradient path and backpropagated through the LLM to update the adversarial suffix.
IV Proposed Approach
@@
We attach a linear probe to each layer $l\in\mathcal{P}$ and optimise a suffix $s$ so that all probes predict the target label $y_{t}\in\{0,1\}$. We define the joint optimisation objective as
$\mathcal{L}_{\text{total}}(s)=\sum_{l\in\mathcal{P}}\lambda_{l}\,\ell^{(l)}\!\left(h^{(l)}(x\oplus s),y_{t}\right),$ (1)
where $x$ denotes the original prompt (including the prompt injection text), $\oplus$ denotes concatenation operator, $\lambda_{l}$ are layer weights, and $\ell^{(l)}$ is the probe loss at layer $l$. In our setting, following [1]each probe is a logistic regression classifier with $z^{(l)}=(W^{(l)})^{\top}h^{(l)}+b^{(l)}$ and $p^{(l)}=\sigma(z^{(l)})$. We initially investigated the use of different values for $\lambda_{l}$ at different layers to put more emphasis on deeper layers. Our intuition was that their activations are more important in detecting prompt injections than earlier layers’ and therefore their gradients must have greater influence in token selection and replacement than the gradients of earlier layers. However, we soon discovered that even $\lambda_{l}\in\{1\}$ can easily fool all the classifiers simultaneously and hence we eschewed weighted gradients.
To integrate probe gradients into the core LLM backpropagation, we backpropagate $\nabla\mathcal{L}_{\text{total}}$ through the LLM and add, at each probed layer, an extra gradient term induced by the probe loss. The resulting gradient recursion for the hidden activations is
@@
$\boxed{\frac{\partial\mathcal{L}_{\text{total}}}{\partial h^{(l)}}=\underbrace{\left(\frac{\partial h^{(l+1)}}{\partial h^{(l)}}\right)^{\top}\frac{\partial\mathcal{L}_{\text{total}}}{\partial h^{(l+1)}}}_{\text{from next layer}}+\underbrace{\mathbf{1}_{\{l\in\mathcal{P}\}}\cdot\lambda_{l}\cdot(p^{(l)}-y_{t})\,W^{(l)}}_{\text{from probe at layer $l$}}}$ (2)
Thus, our attack jointly optimises the adversarial suffix across all probes, while explicitly injecting probe gradients into the LLM gradient flow. Algorithm 1 describes our universal evasive suffix optimisation procedure. The algorithm follows a progressive curriculum strategy: rather than optimising over the full training set at once, which would yield a high-dimensional, ill-conditioned gradient signal, we begin with a single seed prompt and incrementally expand the active set $\mathcal{A}$ as the current suffix achieves sufficient evasion. Specifically, a new prompt is admitted to $\mathcal{A}$ only when the current suffix already fools all probes on at least $\tau_{\mathrm{exp}}$ fraction of the existing active prompts (line 19), ensuring that universality is built up gradually rather than forced prematurely. At each iteration, token-level coordinate gradients are computed for every prompt in $\mathcal{A}$ via the joint backpropagation (LLM layers along with detectors) of Eq. (2), accumulated into a single gradient matrix $G$, and used to sample a pool of candidate suffixes through GCG-style top-$k$ token substitutions (lines 6–12). The candidate minimising the aggregated joint probe loss of Eq. 1 is retained (lines 13–17). Two stopping criteria are enforced: early termination when the suffix universally evades all prompts in the full training set, and a hard iteration budget $K$ (lines 23–28).
Algorithm 1 Universal Evasive Prompt Injection
1:LLM $f$; probes $\{\phi_{l}\}_{l\in\mathcal{P}}$; training prompts $\{x_{i}\}_{i=1}^{N}$; initial suffix $s_{\mathrm{init}}$; target label $y_{t}\in\{0,1\}$; maximum iterations $K$; evasion threshold $\tau\in(0,1)$; batch-expansion threshold $\tau_{\mathrm{exp}}\in(0,1)$
2:Optimised universal suffix $s^{*}$
[…]
5:for $k=1$ to $K$ do
6:  $G\leftarrow\mathbf{0}$ $\triangleright$ Reset accumulated coordinate-grad matrix
7:  $\triangleright$ Step 1: Accumulate gradients over the prompt set
8:  for each $x\in\mathcal{A}$ do
9:   $G_{x}\leftarrow\textsc{TokenGradients}(f,\,x,\,s,\,y_{t})$ $\triangleright$ Joint LLM + probe gradients via Eq. (2)
10:   $G\leftarrow G+G_{x}$ $\triangleright$ Accumulate across prompts to promote universality
11:  end for
12:  $\triangleright$ Step 2: Candidate suffix generation (GCG top-$k$ substitution)
13:  $\{s_{j}\}\leftarrow\textsc{SampleCandidates}(s,\,G)$ $\triangleright$ Draw $B$ candidates by replacing top-$k$ tokens
14:  $\triangleright$ Step 3: Evaluate candidates via the joint probe loss
15:  for each candidate $s_{j}$ do
16:   $\mathcal{L}(s_{j})\leftarrow\displaystyle\sum_{x\in\mathcal{A}}\mathcal{L}_{\mathrm{total}}(x\oplus s_{j})$ $\triangleright$ Eq. (1) summed over the active set
17:  end for
18:  $s\leftarrow\arg\min_{s_{j}}\;\mathcal{L}(s_{j})$ $\triangleright$ Retain the candidate with the lowest joint loss
19:  $\triangleright$ Step 4: Progressive prompt-set expansion
20:  if $\textsc{SuccessRate}(\mathcal{A},\,s)\geq\tau_{\mathrm{exp}}$ then $\triangleright$ Expand only once the current set is mastered
21:   Append $x_{|\mathcal{A}|+1}$ to $\mathcal{A}$ $\triangleright$ Admit the next training prompt in sequence
22:  end if
23:  $\triangleright$ Step 5: Convergence checks
24:  if $\textsc{SuccessRate}(\mathcal{A},\,s)\geq\tau$ and $|\mathcal{A}|=N$ then
25:   break $\triangleright$ All training prompts successfully evaded
26:  end if
27:  if $k=K$ then
28:   break $\triangleright$ Hard iteration budget exhausted
29:  end if
30:end for
31:return $s^{*}\leftarrow s$
V Attack Performace
V-A Setup
@@ V-B Results
V-B1 Attack Success Rate
Table III reports the attack success rate (ASR) under different evaluation metrics. For Phi-3, the suffix achieves an ASR of 93.91% when requiring all five probes to be fooled simultaneously, and 95.28% when considering the last four probes only. Under a majority-vote criterion ($\geq$3 probes), the ASR rises to 99.79%. For Llama-3, the attack is even more effective, reaching 99.63% ASR for both “all five” and “last four” metrics, and nearly perfect success (99.99%) under the majority-vote criterion. These results show that adversarial suffixes can reliably evade probe-based detectors.
@@ V-B2 Distribution of Misclassifications
V-B3 Layer-wise Misclassification
@@
VI Defense
VI-A Approach
We experimented with two approaches to defend against such suffix-based adversarial attacks. The first one is the standard PGD-perturbed [8] adversarial training in which we trained logistic regression models using PGD-perturbed poisoned activations along with the clean and poisoned activations. Since these models have not seen any suffix during their training because no suffix has been used to train them, therefore any optimised suffix that can fool the baseline classifiers can be used to test their robustness against suffix attacks.
The second approach involves training the linear models with the activations obtained by appending suffixes to the prompts. Firstly, we started with different initial suffixes and generated multiple suffixes by targeting the baseline models using the algorithm described in IV and split them into train, validation, and test set. Then, we randomly appended one of the train and validation suffixes to the training and validation prompts respectively and made a forward pass of the LLM to generate activations, which we refer to as adversarially poisoned activations. Finally, we trained logistic regression models with these activations along with the clean and poisoned activations. We tested the robustness of these models by targeting them with test suffixes that the models have not seen before. Furthermore, we also tried to optimise suffixes by directly attacking the adversarially trained models. However, we found that these models are extremely resistant to this type of attack. It’s very difficult to find a suffix despite starting from different initial suffixes and attacking the models directly.
VI-B Defence Performance
Table IV shows the layer-wise performance on clean and poisoned dataset. For clean dataset, suffix trained models’ performance was on par with the baseline models except for the first layer’s classifier whose performance was really poor, while the PGD-trained models performed much worse. For poisoned dataset, PGD-trained models performed best, with the first two classifiers achieving perfect accuracy, while the suffix-trained models’ performance was very close to that of baseline models.
Subset Layer Baseline Models Adv. Trained (PGD) Adv. Trained (Suffix)
Clean 0 71.56% 0% 0.16%